What is the equation line for (6,7);x+8y=9. when y is what?
Answer by richwmiller(17219)

there are two y's
which y do you want
there is the y of (6,7)
and the y of x+8y=9
the given equation line is x+8y=9
6+8*7=9
6+56=9
62=9
62 does not equal 9
so that equation doesn't fit the point (6,7)
So your problem makes no sense.
I could guess what you really are supposed to do but it would be better if you find out what your assignment really is.
And make sure it makes sense!